# Toughness and $A_{\alpha}$-spectral radius in graphs thanks: This work was supported by the Natural Science Foundation of Shandong Province, China (ZR2023MA078)

### v1

@@ Abstract
Let $\alpha\in[0,1)$, and let $G$ be a connected graph of order $n$ with $n\geq f(\alpha)$, where $f(\alpha)=6$ for $\alpha\in[0,\frac{2}{3}]$ and $f(\alpha)=\frac{4}{1-\alpha}$ for $\alpha\in(\frac{2}{3},1)$. A graph $G$ is said to be $t$-tough if $|S|\geq tc(G-S)$ for each subset $S$ of $V(G)$ with $c(G-S)\geq 2$, where $c(G-S)$ is the number of connected components in $G-S$. The $A_{\alpha}$-spectral radius of $G$ is denoted by $\rho_{\alpha}(G)$. In this paper, it is verified that $G$ is a 1-tough graph unless $G=K_{1}\vee(K_{n-2}\cup K_{1})$ if $\rho_{\alpha}(G)\geq\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$, where $\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$ equals the largest root of $x^{3}-((\alpha+1)n+\alpha-3)x^{2}+(\alpha n^{2}+(\alpha^{2}-\alpha-1)n-2\alpha+1)x-\alpha^{2}n^{2}+(3\alpha^{2}-\alpha+1)n-4\alpha^{2}+5\alpha-3=0$. Further, we present an $A_{\alpha}$-spectral radius condition for a graph to be a $t$-tough graph.
@@ 1 Introduction
Let $G$ be an undirected simple graph with vertex set $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$ and edge set $E(G)$. The order of $G$ is the number $n=|V(G)|$ of its vertices and its size is the number $m=e(G)$ of its edges. A graph $G$ is called trivial if $|V(G)|=1$. For arbitrary $v\in V(G)$, the degree of $v$ in $G$ is defined as the number of edges which are adjacent to $v$ and denoted by $d_{G}(v)$. For a given subset $S$ of $V(G)$, we use $G[S]$ to denote the subgraph of $G$ induced by $S$, and write $G-S$ for $G[V(G)\setminus S]$. The complement of a graph $G$ is a graph $\overline{G}$ with the same vertex set as $G$, in which any two distinct vertices are adjacent if and only if they are nonadjacent in $G$. Let $G_{1}$ and $G_{2}$ be two vertex-disjoint graphs. We use $G_{1}\cup G_{2}$ to denote the disjoint union of $G_{1}$ and $G_{2}$. The join $G_{1}\vee G_{2}$ is the graph formed from $G_{1}\cup G_{2}$ by adding all possible edges between $V(G_{1})$ and $V(G_{2})$. We denote by $K_{n}$ a complete graph of order $n$.
Let $t$ be a nonnegative real number. A graph $G$ is said to be $t$-tough if $|S|\geq tc(G-S)$ for each subset $S$ of $V(G)$ with $c(G-S)\geq 2$, where $c(G-S)$ is the number of connected components in $G-S$. The toughness $t(G)$ of $G$ is the largest real number $t$ for which $G$ is $t$-tough, or is $\infty$ if $G$ is complete. This concept was first introduced by Chvátal [1] in 1973. Some results on toughness of graphs can be found in [2, 3, 4, 5, 6].
Given a graph $G$ of order $n$, the adjacency matrix $A(G)$ of $G$ is the $n\times n$ matrix in which entry $a_{ij}$ is 1 or 0 according to whether $v_{i}$ and $v_{j}$ are adjacent or not. The eigenvalues of $G$ are the eigenvalues of its adjacency matrix $A(G)$. Let $\lambda_{1}(G)\geq\lambda_{2}(G)\geq\dots\geq\lambda_{n}(G)$ be its eigenvalues in nonincreasing order. Note that the adjacency spectral radius of $G$ is equal to $\lambda_{1}(G)$, written as $\rho(G)$.
Let $D(G)$ denote the diagonal matrix of vertex degree of $G$. Then $L(G)=D(G)-A(G)$ and $Q(G)=D(G)+A(G)$ are called the Laplacian matrix and signless Laplacian matrix of $G$, respectively. For any $\alpha\in[0,1]$, Nikiforov [7] introduced the $A_{\alpha}$-matrix of $G$ as
@@
Many authors [15, 16, 17, 18, 19, 20] obtained some results on spectral radius in graphs. Brouwer [21] investigated the relationship between toughness and eigenvalues and claimed that for any connected $d$-regular graph $G$, $t(G)>\frac{d}{\lambda}-2$, and he further conjectured that the lower bound $\frac{d}{\lambda}-2$ can be improved to $\frac{d}{\lambda}-1$, where $\lambda=\max_{2\leq i\leq n}|\lambda_{i}|$. Later, Gu [22] strengthened Brouwer’s result and claimed the lower bound $\frac{d}{\lambda}-2$ can be improved to $\frac{d}{\lambda}-\sqrt{2}$. Very recently, Gu [23] confirmed Brouwer’s conjecture. Fan, Lin and Lu [24] presented two adjacency spectral radius conditions for the existence of a $t$-tough graph. It is natural and interesting to find a sufficient condition for a graph to be $t$-tough in view of its $A_{\alpha}$-spectral radius. In this paper, we first show a $A_{\alpha}$-spectral radius condition to guarantee that a graph is 1-tough.
@@
$f(\alpha)=\left\{\begin{array}[]{ll}6,&if\ \alpha\in[0,\frac{2}{3}];\\ \frac{4}{1-\alpha},&if\ \alpha\in(\frac{2}{3},1).\\ \end{array}\right.$
@@
Theorem 1.2. Let $\alpha\in[\frac{1}{2},\frac{3}{4})$, and let $t$ be a positive integer. If $G$ be a connected graph of order $n\geq\max\{4t^{2}+6t+1,\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}\}$ with $\rho_{\alpha}(G)\geq\rho_{\alpha}(K_{2t-1}\vee(K_{n-2t}\cup K_{1}))$, then $G$ is a $t$-tough graph unless $G=K_{2t-1}\vee(K_{n-2t}\cup K_{1})$.
@@ 2 Preliminaries
Proof of Theorem 1.1. Suppose to the contrary that $G$ is not a 1-tough graph, there exists some nonempty subset $S\subseteq V(G)$ such that $c(G-S)\geq|S|+1$. Let $|S|=s$. Then $G$ is a spanning subgraph of $G_{s}^{1}=K_{s}\vee(K_{n_{1}}\cup K_{n_{2}}\cup\cdots\cup K_{n_{s+1}})$ for some integers $n_{1}\geq n_{2}\geq\cdots\geq n_{s+1}$ with $\sum\limits_{i=1}^{s+1}n_{i}=n-s$. According to Lemma 2.2, we have
@@
$-\frac{(2\alpha^{2}-2\alpha)s+\alpha^{2}+\alpha-3}{2(1-\alpha)}<2s+1\leq n,$
@@
If $s=2$, then $-\frac{(2\alpha^{2}-2\alpha)s+\alpha^{2}+\alpha-3}{2(1-\alpha)}=-\frac{5\alpha^{2}-3\alpha-3}{2(1-\alpha)}<6=f(\alpha)\leq n$ due to $0\leq\alpha\leq\frac{2}{3}$. Hence, for $0\leq\alpha\leq\frac{2}{3}$ and $n\geq f(\alpha)=6$, we get
@@
Recall that $\rho_{\alpha}(G_{s}^{2})$ is the largest root of $\varphi_{B_{1}}(x)=0$. As $\eta_{2}<n-2<\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))=\theta(n)$ (see (3.5)), it follows from (3), $\frac{-\alpha^{2}n+2\alpha+s}{2(1-\alpha)}<n-2<\theta(n)$ and $\varphi_{B_{1}}(\theta(n))>0$ that
@@
Proof of Theorem 1.2. Suppose to the contrary that $G$ is not a $t$-tough graph, there exists some nonempty subset $S\subseteq V(G)$ such that $tc(G-S)-1\geq|S|$. Let $|S|=s$ and $c(G-S)=c$. Then $G$ is a spanning subgraph of $G_{1}=K_{tc-1}\vee(K_{n_{1}}\cup K_{n_{2}}\cup\cdots\cup K_{n_{c}})$ for some integers $n_{1}\geq n_{2}\geq\cdots\geq n_{c}$ with $\sum\limits_{i=1}^{c}n_{i}=n-tc+1$. In terms of Lemma 2.2, we conclude
@@
Let $\varphi(c)=(2t+1)c^{2}-(2n+2t+3)c+n^{2}+n+2$. Note that $n\geq(t+1)c-1$. We conclude $3\leq c\leq\frac{n+1}{t+1}$. By a direct computation, it follows from $n\geq\max\{4t^{2}+6t+1,\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}\}$ that
$\varphi(3)-\varphi\left(\frac{n+1}{t+1}\right)=\frac{(n-3t-2)(n-4t^{2}-6t-1)}{(t+1)^{2}}\geq 0,$
which implies that $\varphi(c)$ attains its maximum value at $c=3$ when $3\leq c\leq\frac{n+1}{t+1}$. Combining this with (4), $\alpha\in[\frac{1}{2},\frac{3}{4})$ and $n\geq\max\{4t^{2}+6t+1,\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}\}$, we get
@@
with equality if and only if $G=K_{2t-1}\vee(K_{n-2t}\cup K_{1})$. Which is a contradiction to the $A_{\alpha}$-spectral radius condition of Theorem 1.2. This completes the proof of Theorem 1.2. $\Box$

### v2

@@ Abstract
Let $\alpha\in[0,1)$, and let $G$ be a connected graph of order $n$ with $n\geq f(\alpha)$, where $f(\alpha)=6$ for $\alpha\in[0,\frac{2}{3}]$ and $f(\alpha)=\frac{4}{1-\alpha}$ for $\alpha\in(\frac{2}{3},1)$. A graph $G$ is said to be $t$-tough if $|S|\geq tc(G-S)$ for each subset $S$ of $V(G)$ with $c(G-S)\geq 2$, where c(GS)c(G-S) is the number of connected components in GSG-S. The $A_{\alpha}$-spectral radius of $G$ is denoted by $\rho_{\alpha}(G)$. In this paper, it is verified that $G$ is a 1-tough graph unless $G=K_{1}\vee(K_{n-2}\cup K_{1})$ if $\rho_{\alpha}(G)\geq\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$, where $\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$ equals the largest root of $x^{3}-((\alpha+1)n+\alpha-3)x^{2}+(\alpha n^{2}+(\alpha^{2}-\alpha-1)n-2\alpha+1)x-\alpha^{2}n^{2}+(3\alpha^{2}-\alpha+1)n-4\alpha^{2}+5\alpha-3=0$. Further, we present an $A_{\alpha}$-spectral radius condition for a graph to be a $t$-tough graph.
@@ 1 Introduction
Let $G$ be an undirected simple graph with vertex set $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$ and edge set $E(G)$. The order of $G$ is the number $n=|V(G)|$ of its vertices and its size is the number $m=e(G)$ of its edges. A graph $G$ is called trivial if $|V(G)|=1$. For arbitrary $v\in V(G)$, the degree of $v$ in $G$ is defined as the number of edges which are adjacent to $v$ and denoted by $d_{G}(v)$. For a given subset $S$ of $V(G)$, we use $G[S]$ to denote the subgraph of $G$ induced by $S$, and write GSG-S for $G[V(G)\setminus S]$. The complement of a graph $G$ is a graph $\overline{G}$ with the same vertex set as $G$, in which any two distinct vertices are adjacent if and only if they are nonadjacent in $G$. Let $G_{1}$ and $G_{2}$ be two vertex-disjoint graphs. We use $G_{1}\cup G_{2}$ to denote the disjoint union of $G_{1}$ and $G_{2}$. The join $G_{1}\vee G_{2}$ is the graph formed from $G_{1}\cup G_{2}$ by adding all possible edges between $V(G_{1})$ and $V(G_{2})$. We denote by $K_{n}$ a complete graph of order $n$.
Let $t$ be a nonnegative real number. A graph $G$ is said to be $t$-tough if $|S|\geq tc(G-S)$ for each subset $S$ of $V(G)$ with $c(G-S)\geq 2$, where c(GS)c(G-S) is the number of connected components in GSG-S. The toughness $t(G)$ of $G$ is the largest real number $t$ for which $G$ is $t$-tough, or is $\infty$ if $G$ is complete. This concept was first introduced by Chvátal [6] in 1973. Some results on toughness of graphs can be found in [4, 5, 7, 17, 21, 26, 28, 33, 31, 32].
Given a graph $G$ of order $n$, the adjacency matrix $A(G)$ of $G$ is the $n\times n$ matrix in which entry $a_{ij}$ is 1 or 0 according to whether $v_{i}$ and $v_{j}$ are adjacent or not, and $a_{ij}=0$ if $i=j$. The eigenvalues of $G$ are the eigenvalues of its adjacency matrix $A(G)$. Let $\lambda_{1}(G)\geq\lambda_{2}(G)\geq\dots\geq\lambda_{n}(G)$ be its eigenvalues in nonincreasing order. Note that the adjacency spectral radius of $G$ is equal to $\lambda_{1}(G)$, written as $\rho(G)$.
Let $D(G)$ denote the diagonal matrix of vertex degree of $G$. Then L(G)=D(G)A(G)L(G)=D(G)-A(G) and $Q(G)=D(G)+A(G)$ are called the Laplacian matrix and signless Laplacian matrix of $G$, respectively. For any $\alpha\in[0,1]$, Nikiforov [18] introduced the $A_{\alpha}$-matrix of $G$ as
@@
Many authors [8, 20, 22, 23, 27, 30, 29, 34, 35] obtained some results on spectral radius in graphs. Brouwer [3] investigated the relationship between toughness and eigenvalues and claimed that for any connected $d$-regular graph $G$, $t(G)>\frac{d}{\lambda}-2$, and he further conjectured that the lower bound $\frac{d}{\lambda}-2$ can be improved to $\frac{d}{\lambda}-1$, where $\lambda=\max_{2\leq i\leq n}|\lambda_{i}|$. Later, Gu [11] strengthened Brouwer’s result and claimed the lower bound $\frac{d}{\lambda}-2$ can be improved to $\frac{d}{\lambda}-\sqrt{2}$. Very recently, Gu [10] confirmed Brouwer’s conjecture. Fan, Lin and Lu [9] presented two adjacency spectral radius conditions for the existence of a $t$-tough graph. It is natural and interesting to find a sufficient condition for a graph to be $t$-tough in view of its $A_{\alpha}$-spectral radius. In this paper, we first show a $A_{\alpha}$-spectral radius condition to guarantee that a graph is 1-tough.
@@
[…]
@@
Theorem 1.2. Let $\alpha\in[\frac{1}{2},\frac{3}{4})$, and let $t$ be a positive integer. If $G$ is a connected graph of order $n\geq\max\{5t^{2}+10t+1,\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}\}$ with $\rho_{\alpha}(G)\geq\rho_{\alpha}(K_{2t-1}\vee(K_{n-2t}\cup K_{1}))$, then $G$ is a $t$-tough graph unless $G=K_{2t-1}\vee(K_{n-2t}\cup K_{1})$.
@@ 2 Preliminaries
Proof of Theorem 1.1. Suppose to the contrary that $G$ is not a 1-tough graph, there exists some nonempty subset $S\subseteq V(G)$ such that $c(G-S)\geq|S|+1$. Let $|S|=s$. Then $G$ is a spanning subgraph of $G_{s}^{1}=K_{s}\vee(K_{n_{1}}\cup K_{n_{2}}\cup\cdots\cup K_{n_{s+1}})$ for some positive integers $n_{1}\geq n_{2}\geq\cdots\geq n_{s+1}$ with $\sum\limits_{i=1}^{s+1}n_{i}=n-s$. According to Lemma 2.2, we have
@@
$\frac{(2\alpha^{2}-2\alpha)s+\alpha^{2}+\alpha-3}{2(1-\alpha)}<2s+1\leq n,$
@@
If $s=2$, then $\frac{(2\alpha^{2}-2\alpha)s+\alpha^{2}+\alpha-3}{2(1-\alpha)}=-\frac{5\alpha^{2}-3\alpha-3}{2(1-\alpha)}<6=f(\alpha)\leq n$ due to $0\leq\alpha\leq\frac{2}{3}$. Hence, for $0\leq\alpha\leq\frac{2}{3}$ and $n\geq f(\alpha)=6$, we get
@@
Recall that $\rho_{\alpha}(G_{s}^{2})$ is the largest root of $\varphi_{B_{1}}(x)=0$. From (3), the symmetry axis of $\varphi_{B_{1}}(\theta(n))$ is $\theta(n)=\frac{-\alpha^{2}n+2\alpha+s}{2(1-\alpha)}$. Thus, we see that $\varphi_{B_{1}}(\theta(n))$ is increasing in the interval $\frac{-\alpha^{2}n+2\alpha+s}{2(1-\alpha)},+\infty)$. Note that $\frac{-\alpha^{2}n+2\alpha+s}{2(1-\alpha)}<n-2$. Together with (3.5), we have $\frac{-\alpha^{2}n+2\alpha+s}{2(1-\alpha)}<n-2<\theta(n)$. Combining these with $\varphi_{B_{1}}(\theta(n))>0$, we deduce
@@
Proof of Theorem 1.2. Suppose to the contrary that $G$ is not a $t$-tough graph, there exists some nonempty subset $S\subseteq V(G)$ such that $tc(G-S)-1\geq|S|$. Let $|S|=s$ and c(GS)=cc(G-S)=c, then $tc-1\geq s$. If $n\geq(t+1)c-1$, then $G$ is a spanning subgraph of $G_{1}=K_{tc-1}\vee(K_{n_{1}}\cup K_{n_{2}}\cup\cdots\cup K_{n_{c}})$ for some positive integers $n_{1}\geq n_{2}\geq\cdots\geq n_{c}$ with $\sum\limits_{i=1}^{c}n_{i}=n-tc+1$. In terms of Lemma 2.2, we conclude
@@
Let $\varphi(c)=(2t+1)c^{2}-(2n+2t+3)c+n^{2}+n+2$. Note that $n\geq(t+1)c-1$. We conclude $3\leq c\leq\frac{n+1}{t+1}$. By a direct computation, it follows from $n\geq\max\{5t^{2}+10t+1,\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}\}\geq 5t^{2}+10t+1>4t^{2}+6t+1$ that
$\varphi(3)-\varphi\left(\frac{n+1}{t+1}\right)=\frac{(n-3t-2)(n-4t^{2}-6t-1)}{(t+1)^{2}}>0,$
which implies that $\varphi(c)$ attains its maximum value at $c=3$ when $3\leq c\leq\frac{n+1}{t+1}$. Combining this with (4), $\alpha\in[\frac{1}{2},\frac{3}{4})$ and $n\geq\max\{5t^{2}+10t+1,\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}\}\geq\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}$, we get
@@
with equality if and only if $G=K_{2t-1}\vee(K_{n-2t}\cup K_{1})$. Which is a contradiction to the $A_{\alpha}$-spectral radius condition of Theorem 1.2.
Let $n\leq(t+1)c-2$, then $G$ is a spanning subgraph of $K_{n-c}\vee cK_{1}$ with $c\geq\lceil\frac{n+2}{t+1}\rceil$. Let $G_{3}=K_{n-\lceil\frac{n+2}{t+1}\rceil}\vee\lceil\frac{n+2}{t+1}\rceil K_{1}$. By virtue of Lemma 2.2, we conclude
$\rho_{\alpha}(G)\leq\rho_{\alpha}(K_{n-c}\vee cK_{1})\leq\rho_{\alpha}(G_{3}),$ (4.6)
with equalities if and only if $G=G_{3}$. Note that $2e(G_{3})=(n-\lceil\frac{n+2}{t+1}\rceil)(n-\lceil\frac{n+2}{t+1}\rceil-1)+2\lceil\frac{n+2}{t+1}\rceil(n-\lceil\frac{n+2}{t+1}\rceil)=(n-\lceil\frac{n+2}{t+1}\rceil)(n+\lceil\frac{n+2}{t+1}\rceil-1)<(n-\frac{n+2}{t+1})(n+\frac{n+2}{t+1})$. In view of Lemma 2.5, we possess
$\rho_{\alpha}(G_{3})\leq$ $\frac{2e(G_{3})(1-\alpha)}{n-1}+\alpha n-1$
$<$ $\frac{(1-\alpha)(n-\frac{n+2}{t+1})(n+\frac{n+2}{t+1})}{n-1}+\alpha n-1$
$=$ $n-2+\frac{\psi(n)}{(t+1)^{2}(n-1)},$ (4.7)
where $\psi(n)=-(1-\alpha)n^{2}+((2-\alpha)t^{2}+(4-2\alpha)t+3\alpha-2)n-t^{2}-2t-5+4\alpha$. Notice that
$\frac{(2-\alpha)t^{2}+(4-2\alpha)t+3\alpha-2}{2(1-\alpha)}<5t^{2}+10t+1\leq n$
by $\alpha\in[\frac{1}{2},\frac{3}{4})$, $t\geq 1$ and $n\geq\max\{5t^{2}+10t+1,\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}\}$. For $t\geq 2$, we have
$\psi(n)\leq$ $\psi(5t^{2}+10t+1)$
$=$ $-(1-\alpha)(5t^{2}+10t+1)^{2}+((2-\alpha)t^{2}+(4-2\alpha)t+3\alpha-2)(5t^{2}+10t+1)$
$-t^{2}-2t-5+4\alpha$
$=$ $(5t^{2}+10t+1)((4\alpha-3)t^{2}+(8\alpha-6)t+4\alpha-3)-t^{2}-2t-5+4\alpha$
$\leq$ $(5t^{2}+10t+1)(4(4\alpha-3)+2(8\alpha-6)+4\alpha-3)$
$-t^{2}-2t-5+4\alpha\ \ \ \ \ \ \ \ \ \ \left(\mbox{since}\ t\geq 2\ \mbox{and}\ \alpha<\frac{3}{4}\right)$
$=$ $9(4\alpha-3)(5t^{2}+10t+1)-t^{2}-2t-5+4\alpha$
$<$ $0\ \ \ \ \ \ \ \ \ \ \left(\mbox{since}\ t\geq 2\ \mbox{and}\ \alpha<\frac{3}{4}\right).$
We check that $\psi(n)<0$ also holds for $t=1$. Combining these with (4.5), (4.6) and (4), we conclude
$\rho_{\alpha}(G)\leq\rho_{\alpha}(G_{3})<n-2<\rho_{\alpha}(K_{2t-1}\vee(K_{n-2t}\cup K_{1})),$
which contradicts $\rho_{\alpha}(G)\geq\rho_{\alpha}(K_{2t-1}\vee(K_{n-2t}\cup K_{1}))$. This completes the proof of Theorem 1.2. $\Box$
@@
Acknowledgments
The authors would like to express their sincere gratitude to the anonymous referees for their constructive corrections and valuable comments on this paper, which improved the presentation of this paper. This work was supported by the Natural Science Foundation of Jiangsu Province (Grant No. BK20241949). Project ZR2023MA078 supported by Shandong Provincial Natural Science Foundation.